In a box of switches, it is known that 10% of the switches are faulty. A technician is wiring 30 circuits,



each of which needs one switch. What is the probability that (a) all 30 work (b) at most two of the



switches do not work.



2

1
Expert's answer
2022-02-23T12:26:18-0500

Let the random variable "X" represent the number of faulty switches then, "X\sim Binomial (n=30,p=0.1)" given as,

"p(X=x)=\binom{30}{x}0.1^x0.9^{30-x}, \space x=0,1,2,3,.....,30"


a)

To find the probability that all 30 switches work is equivalent to finding the probability that there are no faulty switches. This can be written as,

"p(X=0)=\binom{30}{0}0.1^00.9^{30}=0.9^{30}=0.04239116"

The probability that all 30 switches work is 0.04239116


b)

"p(X\le 2)=p(x=0)+p(x=1)+p(x=2)=(\binom{30}{0}0.1^00.9^{30})+(\binom{30}{1}0.1^10.9^{29})+(\binom{30}{2}0.1^20.9^{28})=0.04239116+ 0.1413039+ 0.2276562= 0.4113512"

The probability that there are at most 2 faulty switches is 0.4113512